The yield y(t) (in bushes)per acre of a corn crop satisfies the equation dy/dt+y=100+e-t

if y(0)=0 find y at any time t


1
Expert's answer
2021-07-07T09:16:52-0400

Yield is given by, "\frac{dy}{dt}+y=100+e^{-t}"


"I.F. e^{\int dt} =e^t"


Multiplying by Integrating Factor both sides, and integrating

"e^{t}(\frac{dy}{dt}+y)=e^{t}(100+e^{-t})"


"ye^{t} = \int (100 e^{t} + 1) dt"

"ye^{t} = 100 e^{t} + t + C"


Applying condition y(0) = 0,


"0e^{0} = 100 e^{0} + 0 + C \implies C = -100"


"ye^{t} = 100 e^{t} + t -100"


So, yield as a function of t is given by


"y = 100 + t e^{-t}-100e^{-t}"


"y = 100 (1-e^{-t})+ t e^{-t}"